Grade 11General Physics

A circular loog has a radius 10cm and it is carrying a current of 0.1 A.Calculate its magnetic moment.

Khan Tahseen
14 Years agoGrade 11

1 Answer

Aman  Bansal
14 Years ago

Dear Tahseen,

The magnetic moment of circular loop is :

\mathbf{m}=\frac{I}{2}\int\mathbf{r}\times{\rm d}\mathbf{r}.

thus magnetic moment is : pi* 1/2 * 10^-3.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert